OpenStudy (anonymous):

y = x + 3 y = –2x – 7

OpenStudy (aroub):

Solve for x and y?

OpenStudy (anonymous):

yes can you help me

OpenStudy (aroub):

Yup

OpenStudy (aroub):

y=x+3 (Eq1) y=-2x-7 (Eq2) => 2(Eq1)+Eq2 2y=2x+6 y=-2x-7 ---------- 3y=1 y=1/3 1/3=x+3 1=3x+9 -3x=8 x=-8/3 check: y=x+3 1/3=-8/3+3 1/3=1/3

OpenStudy (anonymous):

thank you